OpenStudy (anonymous):

Can someone please explain how to solve this 6y + 2 – 4y = -18

OpenStudy (accessdenied):

We have a few possibilities to start off, although I'd start by combining like terms. Are you familiar with that concept?

OpenStudy (anonymous):

yep

OpenStudy (accessdenied):

Alright, well, we have two terms that have a y, so these are able to be added together first. 6y and -4y

OpenStudy (accessdenied):

6y - 4y = (6 - 4)y = 2y So we can place that into the equation in place of the 6y and -4y: 2y + 2 = -18

OpenStudy (accessdenied):

After that, we simply need to use properties of equality to cancel out everything on the left to isolate y. We are able to add and subtract from both sides, and multiply and divide from both sides as well.

OpenStudy (accessdenied):

So, if we want to cancel out the 2 on the left, we'd subtract 2. But it has to be taken from both sides for equality to hold: \(2y + \cancel{2 \color{red}{- 2}} = -18 \color{red}{-2} \) \(2y + \color{green}0= -20\) \(2y = -20\) Does that make sense?

OpenStudy (accessdenied):

Then similarly for division: \(2y = -20\) \(\displaystyle \frac{\cancel{2}y}{\color{red}{\cancel{2}}} = \frac{-20}{\color{red}{2}} \) and so on. The 2/2 = 1, and we evaluate -20/2 for the solution of y.